# Exploring Equity of Climate Policies using Multi-Agent Multi-Objective Reinforcement Learning

### v1

@@ 1 Introduction
Climate change poses a significant global threat, disproportionately affecting marginalized communities (Faus Onbargi, 2022; Rising et al., 2022). The extent of this threat remains uncertain due to the complex interplay between climate and socioeconomic systems (Burke et al., 2018; van der Wijst et al., 2021). This uncertainty underscores a central challenge of climate justice: the fair allocation of burdens and benefits through the implementation of climate policies (Pozo et al., 2020). The complexity in assessing the implications of policy measures leads to contentious international negotiations, as seen in the Conference of the Parties (COP). Asymmetric impacts, divergent responsibilities, and differing national priorities often result in impasses and heated disagreements among policymakers (Wei et al., 2013; Lamb et al., 2020). Climate change is thus characterized as a wicked problem in the policy domain (Lönngren and Svanström, 2016; Verweij, 2023), necessitating interventions that seek to balance the diverse ethical preferences of stakeholders.
@@
The Intergovernmental Panel on Climate Change (IPCC) is a key player in the global climate change discourse, and its assessment reports serve as the primary source of scientific information for international climate negotiations, such as the COP (Asayama, 2024). These technical reports, crafted by experts from various disciplines, including climate science and economics, provide evidence-based guidance to support effective climate action through the use of Integrated Assessment Models (IAMs) (Cointe et al., 2019). IAMs integrate socioeconomic, climate and technological processes into a single framework to assess global mitigation pathways and simulate future socio-economic and climate scenarios.
@@
Although IAMs are influential tools, they are criticized for suggesting inequitable mitigation policies that disadvantage developing countries (Rivadeneira and Carton, 2022; Gambhir et al., 2022). These models often simplify the complexity of climate change policies to a single objective, which is ethically problematic, as it favours dominant stakeholders and neglects non-economic metrics, such as temperature, biodiversity, and human mortality (Bromley and Beattie, 1973; Kasprzyk et al., 2016). Recent studies seek to address these shortcomings by enhancing IAMs with multiple objectives (Marangoni et al., 2021; Ferrari et al., 2022), or by incorporating multi-agent approaches, such as multi-regional studies (Zhang et al., 2022). However, the multi-objective studies treat the world as a single agent, which overlooks the interests of various stakeholders. Conversely, the multi-agent approach focuses on a single objective, neglecting the diverse preferences of agents during negotiations. Currently, no single modeling framework simultaneously addresses both the multiplicity of objectives and agents, thus overlooking the complexities of the real-world. The exclusion of either multiple objectives or multiple agents limits the applicability of IAMs. Our contribution fills this gap by developing Justice, a Multi-Objective Multi-Agent Reinforcement Learning (MOMARL) IAM framework that emulates real-world negotiations and discovers equitable policy options.  [^1] Code: https://anonymous.4open.science/r/justice-rl-7CC5
@@
MOMARL is a powerful framework for complex decision-making that requires balancing conflicting objectives and coordinating independent decisions over time. In particular, it is a crucial tool for addressing problems that involve sequential decisions (Radulescu, 2024). By extending reinforcement learning (RL) to accommodate multiple agents and multiple objectives, MOMARL enables trade-off management through vectorial rewards, where each component reflects performance on a specific objective. Despite the prevalence of societal problems involving multiple stakeholders with diverse preferences, the field of MOMARL remains underexplored. Simplifications such as hard-coding trade-offs or centralizing decisions limit practical applicability. Importantly, existing works are limited to theoretical examples or simple (grid-like) environments (Felten et al., 2024b; Rădulescu et al., 2020).
@@
Our contribution is twofold. (1) For the IAM community, Justice introduces a multi-objective, multi-agent framework of a climate-economy model, providing a tool that can inform IPCC’s synthesis reports. Such a framework can enable the design of equitable policies that account for regional disparities while balancing multiple aspects of climate change, including economic and environmental outcomes. (2) For the MOMARL community, Justice offers a well-designed, open-source, implementation using the MOMALand API (Felten et al., 2024b), enabling seamless integration with any RL algorithm. This provides a real-world testbed for evaluating and benchmarking future algorithms, supporting the development of robust and societal applications of reinforcement learning methods.
@@ 2.1 Integrated Assessment Models
IAMs offer a holistic approach to inform policy decisions on climate change (Gambhir et al., 2019). These models integrate socioeconomic, technological, and biogeochemical variables to represent interactions between simplified social and climate components (Rivadeneira and Carton, 2022). IAMs primarily inform climate mitigation policies and are frequently used for policy evaluation and optimization (Mastrandrea, 2009). Originating from William Nordhaus’s DICE model [1992], which combined economic and climate models for global cost-benefit analyses of climate policies, IAMs analyze climate policy’s impacts on both the economy and environment. DICE is a foundational framework, earning Nordhaus a Nobel Prize and being utilized by the US government to calculate the social cost of carbon, which measures the societal benefit of reducing CO2 emissions (Grubb et al., 2021). DICE has inspired numerous IAMs and remains a valuable tool for assessing climate policies. A popular regional variation of DICE, known as RICE (Nordhaus and Yang, 1996), encompasses 12 regions, and its recent and enhanced version called RICE50+ (Gazzotti et al., 2021) expands to 57 regions for greater regional resolution.
IAMs can be broadly categorized into two types: (1) highly aggregated Cost-Benefit models and (2) detailed process-based models (Van Beek et al., 2020). Cost-Benefit IAMs (CB-IAMs) are optimization models that identify near-term emission reduction pathways to maximize long-term benefits by considering high-level economic and climate interactions. DICE/RICE IAMs fall under this CB-optimization category. In contrast, Process-Based IAMs (PB-IAMs) are simulation models that offer detailed economic representations across multiple sectors to analyze the impacts of specific policies on economic, social, and environmental factors, with an emphasis on sector-specific environmental impacts, such as those in energy systems (Nikas et al., 2019). CB-IAMs are employed for global mitigation pathways, macroeconomic assessments of mitigation strategies, and strategic interactions reported by the IPCC’s Working Group III (WGIII). Notably, the macroeconomic components of CB-IAMs form the foundation of many process-based models. The global mitigation pathways found by CB-IAMs can also be used to inform the simulation of PB-IAMs. The DICE/RICE family, despite its simplicity, has significantly influenced IPCC WGIII mitigation assessments in various reports, including the recent sixth assessment report, particularly in Chapters 3 and 14, which focus on Mitigation Pathways and International Cooperation, respectively (Shukla et al., 2022).
@@ 2.2 Multi-Objective Multi-Agent RL
We formalize a MOMARL problem as a multi-objective multi-agent Markov decision process (MOMAMDP) with team reward (Rădulescu et al., 2020), defined as a tuple $\bigl{(}\mathcal{S},\mathcal{A},F,\mathbf{R}\bigr{)}$, comprising $N\geq 2$ agents and $d\geq 2$ objectives:
@@
The Pareto front $F(P)$ contains value vectors for Pareto-optimal policies in $P(\boldsymbol{\Pi})$. When the utility function is a positively weighted linear sum, the solution set forms the convex hull of $\mathbf{v}^{\boldsymbol{\pi}}$ (Felten et al., 2024a).
@@ 3.1 High-Level Outline
Justice incorporates the economy, damage, and abatement modules from the RICE50+ model (Gazzotti et al., 2021) and integrates them with the FAIR climate model (Smith et al., 2023), facilitating probabilistic assessments of climate policies across various climate sensitivity scenarios. FAIR is a simplified emulator of complex climate models grounded in climate physics that accurately reproduces historical climate data and is utilized by the IPCC(Leach et al., 2021; Shukla et al., 2022). Unlike other IAMs, Justice identifies Pareto-optimal policies in a multi-objective framework, assessing their robustness across various socioeconomic and climate uncertainties and considering the distributional impacts of the Pareto-optimal policies.
Justice operates on a yearly resolution (instead of a 5-year interval used in RICE50+) and covers 57 independent regions (Table LABEL:tab:justice_macro_regions_grouped in Appendix A [^2] Online Appendix: https://anonymous.4open.science/r/justice-rl-7CC5/Submission_Appendix.pdf). Justice utilizes the Shared Socioeconomic Pathway (SSP) dataset (Riahi et al., 2017) for exogenous regional economic growth, carbon intensity, and population projections, and the Representative Concentration Pathway (RCP) dataset (Meinshausen et al., 2011) informs emission trajectories for other greenhouse gases and aerosols. The SSPs and RCPs provide a consistent framework for integrating socioeconomic and climate research, representing various future scenarios through different socioeconomic narratives and climate forcings. The exogenous SSP data is denoted with a superscript $(^{*})$ in the equations.
Figure 1 provides an overview of the Justice IAM. The economy sub-model employs the neoclassical Cobb-Douglas production function (Gazzotti et al., 2021), as in the DICE IAM, to compute economic growth using exogenous labour and total factor productivity data along with capital stock. The resulting economic output feeds the Emissions sub-model, which calculates CO2 emissions at each timestep based on output and carbon intensity, defined as the fossil fuel share in energy production. The CO2 emissions from various regions are input into the FAIR climate model, which aggregates them to calculate global mean surface temperature, expressed in °C above pre-industrial levels. This global temperature is then downscaled to regional temperature rise using a data-driven statistical downscaler. The downscaled regional temperature is used by the Damage Function to compute the fraction of economic output damaged by temperature increases in every region. Additionally, the Emissions sub-model allows for CO2 mitigation, with associated costs calculated by the Abatement module, based on the emission control rate chosen by agents (or regions) at each timestep. Detailed descriptions and equations for each model component are available in Appendix B 22footnotemark: 2.
@@ 3.2 Multi-Agent Multi-Objective Formulation
We model 12 agents, each representing a macro-region that groups countries following RICE region specification (Nordhaus and Yang, 1996). A map of these macro regions with regional mapping table is provided in Appendix A 22footnotemark: 2. While Justice can generate data for 57 entities, modelling a MOMARL problem with 57 regions is computationally expensive and complicates analysis.
@@ 4 Experimental Setup
We train our agents using Multi-Objective Multi-gent Proximal Policy Optimization (MOMAPPO)(Felten et al., 2024b), an extension of Multi-Agent PPO (MAPPO) (Yu et al., 2022) designed for multi-objective settings. MOMAPPO decomposes a multi-objective problem into multiple single-objective problems using a weighted-sum scalarization function for simplicity. Rewards are normalized to ensure consistency across objectives, and 100 weight vectors are uniformly sampled to generate diverse solutions. For each weight, MOMAPPO trains a multi-agent policy using MAPPO, evaluating its performance and adding non-dominated policies to the solution set. MOMAPPO is trained for 1 million global steps, with evaluations every (approximately) 20,000 steps for 10 random seeds. Appendix C 22footnotemark: 2 includes additional details.
@@ 4.1 Performance Indicators
Evaluating and comparing solution sets in MOMARL is more complex than in single-agent, single-objective RL due to the lack of inherent ordering of solution sets and intertwined performance across objectives. This added complexity leads to varied evaluation methods in MOMARL. To study convergence, we use two most commonly used approaches from both MORL and MARL domains for their suitability in MOMARL settings (Hayes et al., 2022) and to study the equality of distribution between agents we use GINI index:
@@
(Zitzler et al., 2003) represents the region or (hyper-)volume between the points in the solution set and a reference point. The reference point indicates the lower bound for each objective. A solution set can be assessed by comparing its hypervolume with that of competing algorithms or the true Pareto front, if known.
@@
When the decision-maker’s utility function $u$ is linear, the expected utility (EU) metric (Vamplew et al., 2024) can be used to represent the expected utility over a distribution of reward weights $W$.
@@
The GINI index is a measure of inequality that quantifies disparities among agents based on a specific metric, with 0 indicating perfect equality and 1 indicating maximal inequality; we employ Concept-1 GINI by Milanovic (2011) to assess international inequality, where each agent represents a region and the index reflects whether their metrics are converging.
@@ 4.2 RICE50+ Comparison
We compare our results with the default RICE50+ model, as outlined in Section 2.1. This model consists of 57 regions and integrates optimization directly within the global climate simulation. The optimization follows a single-objective approach, employing a social welfare function as the global objective and using a single representative agent to optimize welfare, rather than a multi-agent multi-objective framework. The results presented in the next section are obtained using standard IAM methods, specifically non-linear programming within the General Algebraic Modeling Language (GAMS) for policy optimization. Unlike our approach, RICE50+ does not perform inter-temporal optimization at each timestep. Instead, it optimizes over the entire time horizon by using a social welfare function as a proxy for consumption per capita. Since consumption per capita is directly derived from net economic output (Equation 16a in Appendix B 22footnotemark: 2), we extract the net economic output from RICE50+ to compare with our objectives.
@@ 5.3 Comparison of Key Solutions
trajectories are shown in the center panel of Figure 4. As expected, the best economic policy results in the highest temperature rise, while the best climate policy yields the lowest. Under the compromise policy, temperatures exceed 2°C by the end of the 2100s, reaching about 3.5°C by 2300. The RICE50+ temperature projections are significantly lower than those of the Justice compromise policy over the long term. This discrepancy highlights a common criticism of cost-benefit IAMs: their use of simple climate models inadequately captures the complexities and uncertainties of climate sensitivity, leading to smooth projections that miss abrupt, high-impact events (Mastrandrea and Schneider, 2001; Stanton et al., 2009; Pindyck, 2017). Such models tend to underestimate potential damages and future temperature rises, leading to policy inertia and delayed action (Füssel, 2006; Stern, 2022). We address these issues by using the FAIR model, which captures nonlinear dynamics and feedback mechanisms, effectively handling uncertainty in climate sensitivity (see Appendix B 22footnotemark: 2). FAIR produces temperature scenarios that align with more complex Earth System Models featured in the IPCC (O’Neill et al., 2016).
@@
trajectories are shown in the right panel of Figure 4. All three Justice policies favour rapid near-term mitigation, unlike RICE50+, which delays mitigation and peaks only around 2150. This finding aligns with the criticisms of traditional CB-IAMs for undervaluing early mitigation efforts. Among Justice policies, the best climate policy (purple trajectory) achieves the highest cumulative abated emissions. Both the best economic and compromise policies show similar mitigation levels, though the compromise policy emphasizes earlier action. This finding also aligns with the IPCC’s recommendations for rapid near-term mitigation to limit global warming and promote sustainable growth with minimal climate damage (Shukla et al., 2022).

### v2

@@ 1 Introduction
Climate change poses a significant global threat, disproportionately affecting marginalized communities faus2022climate; rising2022missing. The extent of this threat remains uncertain due to the complex interplay between climate and socioeconomic systems burke2018large; van2021optimality. This uncertainty underscores a central challenge of climate justice: the fair allocation of burdens and benefits through the implementation of climate policies pozo2020reducing. The complexity in assessing the implications of policy measures leads to contentious international negotiations, as seen in the Conference of the Parties (COP). Asymmetric impacts, divergent responsibilities, and differing national priorities often result in impasses and heated disagreements among policymakers wei2013review. Climate change is thus characterized as a wicked problem in the policy domain lonngren2016systems, necessitating interventions that seek to balance the diverse ethical preferences of stakeholders.
@@
The Intergovernmental Panel on Climate Change (IPCC) is a key player in the global climate change discourse, and its assessment reports serve as the primary source of scientific information for international climate negotiations, such as the COP asayama2024history. These technical reports, crafted by experts from various disciplines, including climate science and economics, provide evidence-based guidance to support effective climate action through the use of Integrated Assessment Models (IAMs) cointe2019organising. IAMs integrate socioeconomic, climate and technological processes into a single framework to assess global mitigation pathways and simulate future socio-economic and climate scenarios.
@@
Although IAMs are influential tools, they are criticized for suggesting inequitable mitigation policies that disadvantage developing countries rivadeneira2022justice; gambhir2022climate. These models often simplify the complexity of climate change policies to a single objective, which is ethically problematic, as it favours dominant stakeholders and neglects non-economic metrics, such as temperature, biodiversity, and human mortality bromley1973incongruity; kasprzyk2016battling. Recent studies seek to address these shortcomings by enhancing IAMs with multiple objectives marangoni2021adaptive; ferrari2022optimal, or by incorporating multi-agent approaches, such as multi-regional studies zhang2022ai. However, the multi-objective studies treat the world as a single agent, which overlooks the interests of various stakeholders. Conversely, the multi-agent approach focuses on a single objective, neglecting the diverse preferences of agents during negotiations. Currently, no single modeling framework simultaneously addresses both the multiplicity of objectives and agents, thus overlooking the complexities of the real-world. The exclusion of either multiple objectives or multiple agents limits the applicability of IAMs. Our contribution fills this gap by developing Justice, a Multi-Objective Multi-Agent Reinforcement Learning (MOMARL) IAM framework that emulates real-world negotiations and discovers equitable policy options.  [^1] Code: https://github.com/pollockDeVis/JUSTICE
@@
MOMARL is a powerful framework for complex decision-making that requires balancing conflicting objectives and coordinating independent decisions over time. In particular, it is a crucial tool for addressing problems that involve sequential decisions radulescu2024world. By extending reinforcement learning (RL) to accommodate multiple agents and multiple objectives, MOMARL enables trade-off management through vectorial rewards, where each component reflects performance on a specific objective. Despite the prevalence of societal problems involving multiple stakeholders with diverse preferences, the field of MOMARL remains underexplored. Simplifications such as hard-coding trade-offs or centralizing decisions limit practical applicability. Importantly, existing works are limited to theoretical examples or simple (grid-like) environments felten2024momaland; ruadulescu2020multi.
@@
Our contribution is twofold. (1) For the IAM community, Justice introduces a multi-objective, multi-agent framework of a climate-economy model, providing a tool that can inform IPCC’s synthesis reports. Such a framework can enable the design of equitable policies that account for regional disparities while balancing multiple aspects of climate change, including economic and environmental outcomes. (2) For the MOMARL community, Justice offers a well-designed open-source implementation using the MOMALand API felten2024momaland, allowing seamless integration with any RL algorithm. This provides a real-world testbed for evaluating and benchmarking future algorithms, supporting the development of robust and societal applications of reinforcement learning methods.
@@ 2.1 Integrated Assessment Models
IAMs offer a holistic approach to inform policy decisions on climate change gambhir2019review. These models integrate socioeconomic, technological, and biogeochemical variables to represent interactions between simplified social and climate components rivadeneira2022justice. IAMs primarily inform climate mitigation policies and are frequently used for policy evaluation and optimization mastrandrea2009calculating. Originating from William Nordhaus’s DICE model [nordhaus1992optimal], which combined economic and climate models for global cost-benefit analyses of climate policies, IAMs analyze climate policy’s impacts on both the economy and environment. DICE is a foundational framework, earning Nordhaus a Nobel Prize and being utilized by the US government to calculate the social cost of carbon, which measures the societal benefit of reducing CO2 emissions grubb2021modeling. DICE has inspired numerous IAMs and remains a valuable tool for assessing climate policies. A popular regional variation of DICE, known as RICE nordhaus1996regional, encompasses 12 regions, and its recent and enhanced version called RICE50+ gazzotti2021persistent expands to 57 regions for greater regional resolution.
IAMs can be broadly categorized into two types: (1) highly aggregated Cost-Benefit models and (2) detailed process-based models van2020anticipating. Cost-Benefit IAMs (CB-IAMs) are optimization models that identify near-term emission reduction pathways to maximize long-term benefits by considering high-level economic and climate interactions. DICE/RICE IAMs fall under this CB-optimization category. In contrast, Process-Based IAMs (PB-IAMs) are simulation models that offer detailed economic representations across multiple sectors to analyze the impacts of specific policies on economic, social, and environmental factors, with an emphasis on sector-specific environmental impacts, such as those in energy systems nikas2019detailed. CB-IAMs are employed for global mitigation pathways, macroeconomic assessments of mitigation strategies, and strategic interactions reported by the IPCC’s Working Group III (WGIII). Notably, the macroeconomic components of CB-IAMs form the foundation of many process-based models. The global mitigation pathways found by CB-IAMs can also be used to inform the simulation of PB-IAMs. The DICE/RICE family, despite its simplicity, has significantly influenced IPCC WGIII mitigation assessments in various reports, including the recent sixth assessment report, particularly in Chapters 3 and 14, which focus on Mitigation Pathways and International Cooperation, respectively IPCC2022.
@@ 2.2 Multi-Objective Multi-Agent RL
We formalize a MOMARL problem as a multi-objective multi-agent Markov decision process (MOMAMDP) with team reward ruadulescu2020multi, defined as a tuple $\bigl(\mathcal{S},\mathcal{A},F,\mathbf{R}\bigr)$, comprising $N\geq 2$ agents and $d\geq 2$ objectives:
@@
The Pareto front $F(P)$ contains value vectors for Pareto-optimal policies in $P(\boldsymbol{\Pi})$. When the utility function is a positively weighted linear sum, the solution set forms the convex hull of $\mathbf{v}^{\boldsymbol{\pi}}$ felten2024multi.
@@ 3.1 High-Level Outline
Justice incorporates the economy, damage, and abatement modules from the RICE50+ model gazzotti2021persistent and integrates them with the FAIR climate model smith2023climate, facilitating probabilistic assessments of climate policies across various climate sensitivity scenarios. FAIR is a simplified emulator of complex climate models grounded in climate physics that accurately reproduces historical climate data and is utilized by the IPCC leach2021fairv2; IPCC2022. Unlike other IAMs, Justice identifies Pareto-optimal policies in a multi-objective framework, assessing their robustness across various socioeconomic and climate uncertainties and considering the distributional impacts of the Pareto-optimal policies.
Justice operates on a yearly resolution (instead of a 5-year interval used in RICE50+) and covers 57 independent regions (Table 1 in Appendix A [^2] Online Appendix: https://doi.org/10.5281/zenodo.15424271). Justice utilizes the Shared Socioeconomic Pathway (SSP) dataset sspriahi2017shared for exogenous regional economic growth, carbon intensity, and population projections, and the Representative Concentration Pathway (RCP) dataset meinshausen2011rcp informs emission trajectories for other greenhouse gases and aerosols. The SSPs and RCPs provide a consistent framework for integrating socioeconomic and climate research, representing various future scenarios through different socioeconomic narratives and climate forcings. The exogenous SSP data is denoted with a superscript $(^{*})$ in the equations.
Figure 1 provides an overview of the Justice IAM. The economy sub-model employs the neoclassical Cobb-Douglas production function gazzotti2021persistent, as in the DICE IAM, to compute economic growth using exogenous labour and total factor productivity data along with capital stock. The resulting economic output feeds the Emissions sub-model, which calculates CO2 emissions at each timestep based on output and carbon intensity, defined as the fossil fuel share in energy production. The CO2 emissions from various regions are input into the FAIR climate model, which aggregates them to calculate global mean surface temperature, expressed in °C above pre-industrial levels. This global temperature is then downscaled to regional temperature rise using a data-driven statistical downscaler. The downscaled regional temperature is used by the Damage Function to compute the fraction of economic output damaged by temperature increases in every region. Additionally, the Emissions sub-model allows for CO2 mitigation, with associated costs calculated by the Abatement module, based on the emission control rate chosen by agents (or regions) at each timestep. Detailed descriptions and equations for each model component are available in Appendix B 22footnotemark: 2.
@@ 3.2 Multi-Agent Multi-Objective Formulation
We model 12 agents, each representing a macro-region that groups countries following RICE region specification nordhaus1996regional. A map of these macro regions with regional mapping table is provided in Appendix A 22footnotemark: 2. While Justice can generate data for 57 entities, modelling a MOMARL problem with 57 regions is computationally expensive and complicates analysis.
@@ 4 Experimental Setup
We train our agents using Multi-Objective Multi-gent Proximal Policy Optimization (MOMAPPO)felten2024momaland, an extension of Multi-Agent PPO (MAPPO) yu2022surprising designed for multi-objective settings. MOMAPPO decomposes a multi-objective problem into multiple single-objective problems using a weighted-sum scalarization function for simplicity. Rewards are normalized to ensure consistency across objectives, and 100 weight vectors are uniformly sampled to generate diverse solutions. For each weight, MOMAPPO trains a multi-agent policy using MAPPO, evaluating its performance and adding non-dominated policies to the solution set. MOMAPPO is trained for 1 million global steps, with evaluations every (approximately) 20,000 steps for 10 random seeds. Appendix C22footnotemark: 2 includes additional details.
@@ 4.1 Performance Indicators
Evaluating and comparing solution sets in MOMARL is more complex than in single-agent, single-objective RL due to the lack of inherent ordering of solution sets and intertwined performance across objectives. This added complexity leads to varied evaluation methods in MOMARL. To study convergence, we use two most commonly used approaches from both MORL and MARL domains for their suitability in MOMARL settings hayes2022practical and to study the equality of distribution between agents we use GINI index:
@@
zitzler2003performance represents the region or (hyper-)volume between the points in the solution set and a reference point. The reference point indicates the lower bound for each objective. A solution set can be assessed by comparing its hypervolume with that of competing algorithms or the true Pareto front, if known.
@@
When the decision-maker’s utility function $u$ is linear, the expected utility (EU) metric ruadulescu2020multi can be used to represent the expected utility over a distribution of reward weights $W$.
@@
The GINI index is a measure of inequality that quantifies disparities among agents based on a specific metric, with 0 indicating perfect equality and 1 indicating maximal inequality; we employ Concept-1 GINI by milanovic2011worlds to assess international inequality, where each agent represents a region and the index reflects whether their metrics are converging.
@@ 4.2 RICE50+ Comparison
We compare our results with the default RICE50+ model, as outlined in Section 2.1. This model consists of 57 regions and integrates optimization directly within the global climate simulation. The optimization follows a single-objective approach, employing a social welfare function as the global objective and using a single representative agent to optimize welfare, rather than a multi-agent multi-objective framework. The results presented in the next section are obtained using standard IAM methods, specifically non-linear programming within the General Algebraic Modeling Language (GAMS) for policy optimization. Unlike our approach, RICE50+ does not perform inter-temporal optimization at each timestep. Instead, it optimizes over the entire time horizon by using a social welfare function as a proxy for consumption per capita. Since consumption per capita is directly derived from net economic output (Equation 16a in Appendix B22footnotemark: 2), we extract the net economic output from RICE50+ to compare with our objectives.
@@ 5.3 Comparison of Key Solutions
trajectories are shown in the center panel of Figure 4. As expected, the best economic policy results in the highest temperature rise, while the best climate policy yields the lowest. Under the compromise policy, temperatures exceed 2°C by the end of the 2100s, reaching about 3.5°C by 2300. The RICE50+ temperature projections are significantly lower than those of the Justice compromise policy over the long term. This discrepancy highlights a common criticism of cost-benefit IAMs: their use of simple climate models inadequately captures the complexities and uncertainties of climate sensitivity, leading to smooth projections that miss abrupt, high-impact events mastrandrea2001integrated; stanton2009inside temperature rises, leading to policy inertia and delayed action fussellogical; stern2022time. We address these issues by using the FAIR model, which captures nonlinear dynamics and feedback mechanisms, effectively handling uncertainty in climate sensitivity (see Appendix B22footnotemark: 2). FAIR produces temperature scenarios that align with more complex Earth System Models featured in the IPCC cmipo2016scenario.
@@
trajectories are shown in the right panel of Figure 4. All three Justice policies favour rapid near-term mitigation, unlike RICE50+, which delays mitigation and peaks only around 2150. This finding aligns with the criticisms of traditional CB-IAMs for undervaluing early mitigation efforts. Among Justice policies, the best climate policy (purple trajectory) achieves the highest cumulative abated emissions. Both the best economic and compromise policies show similar mitigation levels, though the compromise policy emphasizes earlier action. This finding also aligns with the IPCC’s recommendations for rapid near-term mitigation to limit global warming and promote sustainable growth with minimal climate damage IPCC2022.